OpenStudy (anonymous):

evaluate this limit for the given value of x and function f: f(x)= 1/x and x=-2 in the limit of the form: f(x+h)- f(x)/h lim h-->0

OpenStudy (nikita2):

it is just derivative f'(-2) So lim = -(1/(-2)^2)= -1/4

OpenStudy (anonymous):

I don't understand:-(

OpenStudy (nikita2):

Do you know what is derivative?

OpenStudy (anonymous):

Nope. Just started calc; this is first hw question on limits...

OpenStudy (nikita2):

Soz)

OpenStudy (anonymous):

I appreciate your help!

OpenStudy (nikita2):

lim (1/(-2+h) - 1/(-2))/h = lim ( (-2 - (-2 +h) )/(-2+h)*(-2) )/h = lim (-(h )/(-2+h)*(-2) )/h = lim -1/ (-2+h)*(-2) = -1/4

OpenStudy (anonymous):

AH! Thank you!